OpenStudy (anonymous):

how do you find the area of an equilateral triangle if it has 14 inch sides?????????

OpenStudy (a1234):

A = 1/2bh 1/2(14) = 7 (7)(14) = 98 I think this is how you do it, but I'm not sure.

OpenStudy (anonymous):

If all they give you is the length of the sides, you cant use the 1/2bh formula because you dont have the height. Instead you can use the formula sqrt(3)/4(l)^2 where l=length (14 in this case)

OpenStudy (a1234):

Oh, okay. Sorry, I actually haven't done such problems.

OpenStudy (anonymous):

thanks guys for the help! @steffw and @a1234